OpenStudy (anonymous):

log 0.06 How do you do it NOT on a calculator?

OpenStudy (anonymous):

r u given anything else? for example are you given log(6) = ???

OpenStudy (anonymous):

nope. it a homework problem and all it says is log 0.06 :/

OpenStudy (anonymous):

in that case then, you just need to use the properties of logs: \(\large log(0.06)=log(\frac{6}{100})=log(6)-log(100)=log(6)-2 \) that is the best you can do without a calculator.

OpenStudy (anonymous):

alright, ill put that and see what my teach says, thank you!

OpenStudy (anonymous):

yw...